# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v4

@@ 4.3 Do existing inversion methods fix the correlations?
Finally, we investigate whether inversion methods introduced in prior work resolve the issue of selectively reduced latents diversity. To this end, we employ the Stable Diffusion XL (Podell et al., 2024) model, and using $2000$ generations from COCO-30K prompts, we measure resulting inversions’ normality, editability, and image reconstruction performance. Results in Table 7 indicate that methods based on predicted noise regularization, such as Pix2Pix-Zero (Parmar et al., 2023) and ReNoise (Garibi et al., 2024), while slightly improving the latents’ quality, do not offer significantly better editability, while being two times slower than DDIM. On the other hand, replacing DDIM approximation (Eq. 3) with reverse DPMSolver (Lu et al., 2022) leads to decorellated latents at the cost of high image reconstruction error. We show that our fix offers the best editability of latents with minimal reconstruction loss, all at the lowest computational cost. To be more precise, thanks to the fact that selected inversions steps replaced with the randomly sampled noise are the least important in terms of accurate reconstruction, we can observe no increase in the reconstruction error when replacing 2% of forward steps (up to the 2nd decimal point of MAE), while for 4% of steps the additional error is around 1% of pixel deviations - a value below the threshold usually employed by adversarial attacks as being not noticeable by a human eye (Madry et al., 2017).
@@
Prior NFEs Normality Image Reconstruction CLIP Text Alignment Inv. time
@@
Corr. $\downarrow$ KL $\times 10^{-2}\downarrow$ MAE $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ Source $\downarrow$ Target $\uparrow$ [s/image] $\downarrow$
Gaussian Noise $0.08_{\pm.01}$ 0.10 $31.88_{\pm 11.66}$ $73.34_{\pm 9.65}$
DDIM Inv. 50 $0.16_{\pm.02}$ 0.89 0.03 0.10±.05 27.58 $34.99_{\pm 11.36}$ $69.58_{\pm 10.17}$ $7.17_{\pm.01}$
Pix2Pix-Zero 50 $0.15_{\pm.02}$ 0.85 0.03 0.10±.05 27.35 $34.86_{\pm 11.39}$ $69.73_{\pm 10.12}$ $22.07_{\pm 1.84}$
ReNoise ($T=50,K=1$) 50 $0.14_{\pm.02}$ 0.73 0.04 0.09±.05 25.64 $34.89_{\pm 11.46}$ $69.87_{\pm 10.07}$ $19.86_{\pm.22}$
ReNoise ($T=25,K=2$) 50 $0.14_{\pm.02}$ 0.59 0.04 0.09±.04 24.81 $35.21_{\pm 11.61}$ $69.68_{\pm 10.09}$ $15.36_{\pm.51}$
ReNoise ($T=17,K=3$) 51 0.13±.02 0.47 0.06 $0.13_{\pm.05}$ 22.35 $35.79_{\pm 11.65}$ $69.04_{\pm 9.98}$ $14.31_{\pm.49}$
DPMSolver-1 ($T=50$) 50 0.09±.01 0.50 0.06 $0.30_{\pm.10}$ 22.55 $34.81_{\pm 11.40}$ $70.26_{\pm 10.42}$ $7.06_{\pm.00}$
DPMSolver-2 ($T=25$) 50 0.09±.01 0.26 0.06 $0.14_{\pm.07}$ 24.76 $34.69_{\pm 11.55}$ 71.24±9.91 $7.06_{\pm.00}$
Ours (forward 2%) 49 $0.14_{\pm.02}$ 0.71 0.03 0.10±.05 27.12 34.32±11.49 $70.21_{\pm 10.13}$ 7.00±.00
Ours (forward 4%) 48 0.09±.01 0.38 0.04 $0.14_{\pm.04}$ 25.68 $\mathbf{33.62}_{\pm 11.63}$ $\mathbf{72.17}_{\pm 9.94}$ 6.86±.01
@@ 5 Conclusions
In the Appendix we first outline the limitations (A) of our experiments, discuss the broader impact (B) of the proposed fix, usage of LLMs (C), and list the computational resources (D) used. We then describe, in detail, the DDIM approximation error (E) and the consequences it has across varying numbers of sampling steps (F). Next, we describe our experiments: how we condition DMs (G), measuring the noise–image–latent triangles (H), identifying plain-regions pixels (I) in the image, and computing the inversion error (J). In section K, we provide the full pseudocode for the proposed fix. We demonstrate how the technique enhances image diversity and prompt alignment (L) and how it enables stochastic editing of real images (M). Finally, we compare Gaussian noise and DDIM latents in their mappings to images (N) and track how these relationships evolve during DM training (O). In section P, we include additional qualitative results: when the introduced fix might fail, and how it enhances image interpolation.
@@ Appendix A Limitations
In this work, we analyzed the relation between the random noise, image generations and their latent encodings obtained through DDIM Inversion. While our studies focused on DDIM approximation error from Song et al. (2021), there exist other solvers and inversion methods, as described in Section 2, bringing their own advantages and limitations. The error of DDIM inversion strongly depends on the number of steps with which it is performed. In particular, performing the process very granularly, e.g., using $T=1000$ steps, can result in strong suppression of the correlation. Nevertheless, the default number of steps we have chosen, i.e., 100, is, according to previous works (Hong et al., 2024; Garibi et al., 2024; Kim et al., 2022b), a practical choice as a good balance between the reconstruction error and the speed of the algorithm. In Appendix F, we present that the latents exhibit correlations when using $1000$ sampling steps, and that the proposed fix can help also in such case.
@@
As mentioned in Section 2, the DDIM inversion error can be detrimental when using a small number of steps. Even though the solution proposed in this work (involving forward diffusion in first inversion steps) drastically removes correlations in latents and, thus, improves image interpolation and editing, it does not improve the numerical inversion error resulting from using small number of steps. In our experiments with 50 steps that are commonly used for edition, we show no significant drawbacks. However, in the extreme cases, using our fix in even a single step, might result in the loss of information necessary for correct image reconstruction, hence it may be then less preferred than standard DDIM inversion. In Section P.1, we present failure cases for introduced solution.
@@ Appendix C Broader impact
As our work is mostly analytical, we do not provide new technologies that might have a significant societal impact. However, our solution for improving the accuracy of DDIM inversion has potential implications that extend beyond technical advancements in diffusion models. As our fix enables more prompt-aligned image editing it could be misused to advance image manipulation techniques. The enhanced interpolation quality could make synthetic content more convincing and harder to detect. The authors do not endorse using the method for deceptive or malicious purposes, and discourage any application that could erode trust or cause harm.
@@ Appendix D Compute resources
Almost all the experiments we perform are based on performing a sampling process using a diffusion model from noise, performing DDIM inversion and, possibly, image reconstruction from the latent, where each of these processes takes the same number of steps, hence the same number of GPU-hours on average. As our experiments differ in terms of number of sampling steps and images to generate, we provide average GPU time per one sampling step per batch (with $B$ denoting batch size) for each model as following: ADM-32 ($B=256$): 0.054s, ADM-64 ($B=128$): 0.093s, ADM-256 ($B=64$): 0.901s, LDM ($B=128$): 0.273s, DiT ($B=128$): 0.104s, IF ($B=64$): 0.609s. Note that some experiments, such as analyzing inversion approximation errors per step (Fig. 4) or sampling from noise interpolations (Fig. 5), involves performing the procedures several times. Taking into consideration all the experiments described in the main text of this work, fully reproducing them takes roughly 110 GPU hours. However, considering the prototyping time, preliminary and failed experiments, as well as the fact that most of the experiments must be performed sequentially (e.g., inversion after image generation, image reconstruction after inversion), the overall execution time of the entire research project is multiple times longer.
@@ Appendix E Approximation error in DDIM Inversion
While setting $\eta=1$ makes Eq. 7 equivalent to Eq. 6, leading to a Markovian probabilistic diffusion model, setting $\eta=0$ removes the random component from the equation, making it a Denoising Diffusion Implicit Model (DDIM), which prominent ability is to perform a deterministic mapping from given noise ($x_{T}$) to image ($x_{0}$). One of the potential benefits of implicit models is the possibility of reversing the backward diffusion process to transfer images back to the original noise. Operating in such a space by modifying resulting inversions unlocks numerous image manipulation capabilities, i.a., image editing (Hertz et al., 2022; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Parmar et al., 2023; Rout et al., 2025; Miyake et al., 2023; Brack et al., 2024; Tang et al., 2024; Hong et al., 2024; Wallace et al., 2023; Samuel et al., ; Garibi et al., 2024; Pan et al., 2023; Dong et al., 2023), image interpolation (Zheng et al., 2024; Zhang et al., 2024b; Samuel et al., 2023; Dhariwal & Nichol, 2021), or stroke-to-image synthesis (Meng et al., 2022; Rout et al., 2025). The inversion process can be derived from Eq. 7, leading to the formula for transferring a less noisy image $x_{t-1}$ to a more noisy one $x_{t}$:
@@
Unfortunately, a perfect image-to-noise inversion is infeasible. Due to circular dependency on $x_{t}$ within Eq. 8, Dhariwal & Nichol (2021) propose to approximate this equation by assuming that the model’s prediction in $t$-th step is locally equivalent to the decision in (t1)(t-1)-th step $\epsilon_{\theta}^{(t)}(x_{t})\approx\epsilon_{\theta}^{(t)}(x_{t-1})$. The inverted trajectory is determined in multiple steps. Hence, the error propagates further away from the image, leading to the latents that deviate from the starting noise. This flaw is presented in Fig. 8.
@@
Figure 8: The DDIM inversion error, derived from approximating DM’s prediction in step $t$ with the $(t+1)$ output (left), propagates with the next inversion steps, leading to a distribution of latents $\hat{x}^{T}$ that deviates significantly from the expected noise distribution $x^{T}$ (right).
Appendix F Correlations vs. inversion steps

### v5

@@ 4.3 Do existing inversion methods fix the correlations?
So far, we have demonstrated the issues of the classical DDIM inversion method. In this section, we investigate whether novel inversion methods introduced in prior work resolve the issue of selectively reduced latents’ diversity. We employ the Stable Diffusion XL (Podell et al., 2024) model, and using $2000$ generations from COCO-30K prompts, we measure resulting inversions’ normality, editability, and image reconstruction performance. For fair comparison, we use the same number of NFEs. Results in Table 7 indicate that methods based on predicted noise regularization, such as Pix2Pix-Zero (Parmar et al., 2023) and ReNoise (Garibi et al., 2024), while slightly improving the latents’ quality, do not offer significantly better editability, while being two times slower than DDIM. On the other hand, replacing DDIM approximation (Eq. 3) with reverse DPMSolver (Lu et al., 2022) leads to decorellated latents at the cost of high image reconstruction error. We show that our fix offers the best editability of latents with minimal reconstruction loss, all at the lowest computational cost. To be more precise, thanks to the fact that selected inversions steps replaced with the randomly sampled noise are the least important in terms of accurate reconstruction, we can observe no increase in the reconstruction error when replacing 2% of forward steps (up to the 2nd decimal point of MAE), while for 4% of steps the additional error is around 1% of pixel deviations - a value below the threshold usually employed by adversarial attacks as being not noticeable by a human eye (Madry et al., 2017). This replacement percentage allows us to navigate the trade-off between reconstruction fidelity and latent manipulability.
@@
Prior NFE Normality Image Reconstruction CLIP Text Alignment Inv. time
@@
Corr. $\downarrow$ KL $\downarrow$ MAE $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ Source $\downarrow$ Target $\uparrow$ Direct. $\uparrow$ [s/image] $\downarrow$
Gaussian Noise $0.08_{\pm.01}$ 0.10 $31.88_{\pm 11.66}$ $73.34_{\pm 9.65}$ $80.62_{\pm 16.96}$
DDIM Inv. 50 $0.16_{\pm.02}$ 0.89 0.03 0.10±.05 27.58 $34.99_{\pm 11.36}$ $69.58_{\pm 10.17}$ $75.59_{\pm 17.86}$ $7.17_{\pm.01}$
Pix2Pix-Zero 50 $0.15_{\pm.02}$ 0.85 0.03 0.10±.05 27.35 $34.86_{\pm 11.39}$ $69.73_{\pm 10.12}$ $75.83_{\pm 17.87}$ $22.07_{\pm 1.84}$
ReNoise (T=50, K=1) 50 $0.14_{\pm.02}$ 0.73 0.04 0.09±.05 25.64 $34.89_{\pm 11.46}$ $69.87_{\pm 10.07}$ $76.47_{\pm 18.12}$ $19.86_{\pm.22}$
ReNoise (T=25, K=2) 50 $0.14_{\pm.02}$ 0.59 0.04 0.09±.04 24.81 $35.21_{\pm 11.61}$ $69.68_{\pm 10.09}$ $76.17_{\pm 18.15}$ $15.36_{\pm.51}$
ReNoise (T=17, K=3) 51 0.13±.02 0.47 0.06 $0.13_{\pm.05}$ 22.35 $35.79_{\pm 11.65}$ $69.04_{\pm 9.98}$ $75.20_{\pm 17.96}$ $14.31_{\pm.49}$
DPMSolver-1 (T=50) 50 0.09±.01 0.50 0.06 $0.30_{\pm.10}$ 22.55 $34.81_{\pm 11.40}$ $70.26_{\pm 10.42}$ $74.76_{\pm 18.02}$ $7.06_{\pm.00}$
DPMSolver-2 (T=25) 50 0.09±.01 0.26 0.06 $0.14_{\pm.07}$ 24.76 $34.69_{\pm 11.55}$ 71.24±9.91 $76.17_{\pm 18.10}$ $7.06_{\pm.00}$
Ours (forward 2%) 49 $0.14_{\pm.02}$ 0.71 0.03 0.10±.05 27.12 34.32±11.49 $70.21_{\pm 10.13}$ 76.76±17.79 7.00±.00
Ours (forward 4%) 48 0.09±.01 0.38 0.04 $0.14_{\pm.04}$ 25.68 $\mathbf{33.62}_{\pm 11.63}$ $\mathbf{72.17}_{\pm 9.94}$ 78.91±17.49 6.86±.01
@@
4.4 Improving state-of-the-art editing engines with our fix
Finally, we evaluate the possibility of combining our simple fix with existing methods designed for real image manipulation. We adapt StyleAligned (Hertz et al., 2024), the state-of-the-art method for transferring a style from a reference image to new generations, and MasaCtrl (Cao et al., 2023), a complex editing engine for text-based real image editing. As these methods employ Naïve DDIM Inversion to find starting noise for input images, we can directly apply our simple fix to those techniques, without changing their generation procedure.
Inversion Method
CLIP Prompt
Alignment $\uparrow$
Set Consistency
(DINO) $\uparrow$
Set Consistency
(CSD) $\uparrow$
Style Similarity
(DINO) $\uparrow$
Style Similarity
(CSD) $\uparrow$
Naive DDIM 0.795 0.476 0.552 0.505 0.690
Ours (forward 4%) 0.795 0.471 0.554 0.510 0.697
Table 8: Style transfer from reference image with StyleAligned (Hertz et al., 2024) incorporating Naïve DDIM Inversion and version with our fix. Our fix improves similarity to input style.
We evaluate style transfer by measuring generations’ alignment to the prompt, set consistency (pairwise cosine similarities of DINO (Darcet et al., 2024) and CSD (Somepalli et al., 2025) embeddings), and style consistency to the reference image (DINO and CSD embeddings cosine similarity). The Table 8 compares the performance of vanilla StyleAligned and the version with our fix in style transfer from StyleDrop (Sohn et al., 2023) images. As presented, our approach improves the alignment with the target style. Additionally, in Fig. 8, we present a qualitative comparison of both inversion algorithms when combined with StyleAligned (1) for style transfer and MasaCtrl (2) for real image editing. More examples can be found in Sections O.4 and O.5.
Figure 8: DDIM Inversion with our fix, when merged to popular image manipulation engines, improves (1) style transfer with StyleAligned and (2) image editing with MasaCtrl. We use real images from, accordingly, StyleDrop (Sohn et al., 2023) and PIEBench (Ju et al., 2024) benchmarks.
@@ 5 Conclusions
In the Appendix, we first outline the limitations (A) of our experiments, LLM usage during writing (B), discuss the broader impact (C) of this work, and list the computational resources (D) used. Following, we describe, in detail, the DDIM approximation error (E), and the fix we introduce (pseudocode) in this work (F). Next, we describe our experiments: measuring the noise–image–latent triangles (G), methodology for identifying plain-regions pixels (H) in the image, and computing the inversion error (I). In (J), we demonstrate how we condition the models, and in (K), we present more details on image diversity and prompt alignment during editing. Next, we compare Gaussian noise and DDIM latents in their mappings to images (L) and track how these relationships evolve during DM training (M). In section (N) we discuss the impact of different parameters’ values: number of inversion steps used, number of inversion steps replaced with forward diffusion, and the impact of guidance scale. In (O), we include additional qualitative results for image interpolation, reconstruction, editing, and style transfer. Finally, in section (P), we present that the investigated issue with latents’ correlations also exists in Flow Matching models.
@@ Appendix A Limitations
In this work, we analyzed the relation between the random noise, image generations and their latent encodings obtained through DDIM Inversion. While our studies focused on DDIM approximation error from Song et al. (2021), there exist other solvers and inversion methods, as described in Section 2, bringing their own advantages and limitations. The error of DDIM inversion strongly depends on the number of steps with which it is performed. In particular, performing the process very granularly, e.g., using $T=1000$ steps, can result in strong suppression of the correlation. Nevertheless, the default number of steps we have chosen, i.e., 100, is, according to previous works (Hong et al., 2024; Garibi et al., 2024; Kim et al., 2022b), a practical choice as a good balance between the reconstruction error and the speed of the algorithm. In Section N.1, we present that the latents exhibit correlations when using $1000$ sampling steps, and that the proposed fix can also help in such cases.
@@
As mentioned in Section 2, the DDIM inversion error can be detrimental when using a small number of steps. Even though the solution proposed in this work (involving forward diffusion in first inversion steps) drastically removes correlations in latents and, thus, improves image interpolation and editing, it does not improve the numerical inversion error resulting from using a small number of steps. In our experiments with 50 steps that are commonly used for editing, we show no significant drawbacks. However, in the extreme cases, using our fix in even a single step might result in the loss of information necessary for correct image reconstruction, hence it may then be less preferred than standard DDIM inversion. In Section N.1, we present failure cases for the introduced solution.
@@ Appendix C Broader impact
As our work is mostly analytical, we do not provide new technologies that might have a significant societal impact. However, our solution for improving the accuracy of DDIM inversion has potential implications that extend beyond technical advancements in diffusion models. As our fix enables more prompt-aligned image editing, it could be combined with various editing engines and misused to advance image manipulation techniques. The enhanced interpolation quality could make synthetic content more convincing and harder to detect. The authors do not endorse using the method for deceptive or malicious purposes, and discourage any application that could erode trust or cause harm.
@@ Appendix D Compute resources
Almost all the experiments we perform are based on performing a sampling process using a diffusion model from noise, performing DDIM inversion, and, possibly, image reconstruction from the latent, where each of these processes takes the same number of steps, hence the same number of GPU-hours on average. As our experiments differ in terms of number of sampling steps and images to generate, we provide average GPU time per one sampling step per batch (with $B$ denoting batch size) for each model as following: ADM-32 ($B=256$): 0.054s, ADM-64 ($B=128$): 0.093s, ADM-256 ($B=64$): 0.901s, LDM ($B=128$): 0.273s, DiT ($B=128$): 0.104s, IF ($B=64$): 0.609s. Note that some experiments, such as analyzing inversion approximation errors per step (Fig. 4) or sampling from noise interpolations (Fig. 5), involves performing the procedures several times. Taking into consideration all the experiments described in the main text of this work, fully reproducing them takes roughly 110 GPU hours. However, considering the prototyping time, preliminary and failed experiments, as well as the fact that most of the experiments must be performed sequentially (e.g., inversion after image generation, image reconstruction after inversion), the overall execution time of the entire research project is multiple times longer.
@@ Appendix E Approximation error in DDIM Inversion
While setting $\eta=1$ makes Eq. 7 equivalent to Eq. 6, leading to a Markovian probabilistic diffusion model, setting $\eta=0$ removes the random component from the equation, making it a Denoising Diffusion Implicit Model (DDIM), which prominent ability is to perform a deterministic mapping from given noise ($x_{T}$) to image ($x_{0}$). One of the potential benefits of implicit models is the possibility of reversing the backward diffusion process to transfer images back to the original noise. Operating in such a space by modifying resulting inversions unlocks numerous image manipulation capabilities, i.a., image editing (Hertz et al., 2022; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Parmar et al., 2023; Rout et al., 2025; Miyake et al., 2023; Brack et al., 2024; Tang et al., 2024; Hong et al., 2024; Wallace et al., 2023; Samuel et al., 2025; Garibi et al., 2024; Pan et al., 2023; Dong et al., 2023), image interpolation (Zheng et al., 2024; Zhang et al., 2024b; Samuel et al., 2023; Dhariwal and Nichol, 2021), or stroke-to-image synthesis (Meng et al., 2022; Rout et al., 2025). The inversion process can be derived from Eq. 7, leading to the formula for transferring a less noisy image $x_{t-1}$ to a more noisy one $x_{t}$:
@@
Unfortunately, a perfect image-to-noise inversion is infeasible. Due to circular dependency on $x_{t}$ within Eq. 8, Dhariwal and Nichol (2021) propose to approximate this equation by assuming that the model’s prediction in $t$-th step for $x_{t}$ is locally equivalent to the decision for $x_{t-1}$: $\epsilon_{\theta}^{(t)}(x_{t})\approx\epsilon_{\theta}^{(t)}(x_{t-1})$. The inverted trajectory is determined in multiple steps. Hence, the error propagates further away from the image, leading to the latents that significantly deviate from the starting noise. This flaw is presented in Fig. 9.
@@
Figure 9: The DDIM inversion error, derived from approximating DM’s prediction for $x_{t}$ with the output for $x_{t-1}$ (left), propagates with the next inversion steps, leading to a distribution of latents $\hat{x}^{T}$ that deviates significantly from the expected noise distribution $x^{T}$ (right).
Appendix F Pseudocode for DDIM Inversion with Forward Diffusion
We present in Algorithm 1 the pseudocode for the proposed solution to the decorrelation of latent encodings by replacing the first $t^{\prime}$ inversion steps with a forward diffusion process. In Sections N.1, N.2 and N.3, we analyze the sensitivity of this fix – how it performs for: different number of inversion steps $T$ (Section N.1), different percentage of inversion steps replaced with forward diffusion $f$ (Section N.2), and when classifier-free guidance is applied (Section N.3).
2:decorrelated latent encoding $\mathbf{\hat{x}_{T}}$}
3:sample $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$
[…]
5:for $t\leftarrow f+1,\dots,T$ do
7:end for
8:return $\mathbf{\hat{x}_{T}}$}
Appendix G Most probable triangles
G.1 Methodology
G.2 Results
Model $T$ $\angle{\mathbf{x_{0}}}$ $\angle{\mathbf{x_{T}}}$ $\angle{\mathbf{\hat{x}_{T}}}$
ADM $32\times 32$ $10$ $44$ $16$ $120$
$100$ $29$ $28$ $123$
$1000$ 2020 $45$ $115$
ADM $64\times 64$ $10$ $30$ $31$ $119$
$100$ $11$ $60$ $109$
$1000$ 66 $79$ $95$
ADM $256\times 256$ $10$ $24$ $50$ $106$
$100$ $24$ $73$ $83$
$1000$ $23$ $73$ $84$
LDM $256\times 256$ $10$ $23$ $53$ $104$
$100$ 22 $76$ $102$
$1000$ 11 $83$ $96$
DiT $256\times 256$ $10$ $27$ $47$ $106$
$100$ 44 $66$ $110$
$1000$ 11 $80$ $99$
(a) ADM-32 (CIFAR-10)
(b) ADM-256 (ImageNet)
(c) LDM (CelebA)
G.3 Example histograms
Appendix H Plain surface thresholding
During the experiments, we set $\tau=0.025$.
$\mathcal{M}(h,w)=\prod_{c}\mathcal{M}_{c}(h,w).$ (12)
Appendix I Calculating inversion error
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in Appendix H. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
and adequately for non-plain ($n$) pixels as
Pixel area Diffusion steps Model
ADM-64 DiT
Plain $1,2\dots 50$ $100\%$ $15.11$ $5.67$
Non-plain $12.43$ $3.16$
Plain $1,2\dots 5$ $10\%$ $9.23$ $3.41$
Non-plain $6.80$ $1.42$
Plain $6,7\dots 50$ $90\%$ $5.87$ $2.23$
Non-plain $5.63$ $1.74$
Appendix J Conditioning for Diffusion Models
For a thorough analysis, we employ both unconditional (ADM-32, ADM-64, ADM-256) and conditional (DiT, IF, SDXL) diffusion models. For conditioning, we take prompts from the Recap-COCO-30K dataset (Li et al., 2025) for IF and Stable Diffusion XL and ImageNet class names for DiT. However, as noted by Mokady et al. (2023), Classifier-Free Guidance introduces additional errors to the DDIM inversion. To focus solely on the inversion approximation error, in the experiments from the main part of this work, we disable CFG by setting the guidance scale to $w=1$. However, in Section N.3, we show that the proposed fix can also improve DDIM Inversion when CFG is enabled ($w>1$).
Appendix K Details on diversity and alignment of edition
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
Noise $\mathbf{x_{T}}$} $0.709$ $0.591$ $0.228$ $0.174$ $0.465$ $0.681$
DDIM Latent $\mathbf{\hat{x}_{T}}$} $0.677$ $0.564$ $0.261$ $0.223$ $0.480$ $0.662$
11 ($2\%$) $0.696$ $0.580$ $0.239$ $0.187$ $0.470$ $0.676$
$1\dots 2$ ($4\%$) $0.701$ $0.587$ $0.227$ $0.179$ $0.469$ $0.677$
$1\dots 3$ ($6\%$) $0.705$ $0.594$ $0.216$ $0.173$ $0.468$ $0.678$
$1\dots 5$ ($10\%$) $0.711$ $0.605$ $0.198$ $0.165$ $0.466$ $0.679$
$1\dots 50$ ($100\%$) $0.805$ $0.801$ $0.013$ $0.085$ $0.465$ $0.680$
(a) Diffusion Transformer
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
Noise ($100\%$) $0.665$ $0.380$ $0.339$ $0.344$ $0.273$ $0.649$
DDIM Latent $\mathbf{\hat{x}_{T}}$} $0.609$ $0.328$ $0.406$ $0.416$ $0.353$ $0.614$
11 ($2\%$) $0.666$ $0.376$ $0.359$ $0.335$ $0.281$ $0.646$
$1\dots 2$ ($4\%$) $0.660$ $0.369$ $0.359$ $0.351$ $0.287$ $0.649$
$1\dots 3$ ($6\%$) $0.661$ $0.370$ $0.353$ $0.349$ $0.285$ $0.650$
$1\dots 5$ ($10\%$) $0.662$ $0.372$ $0.341$ $0.346$ $0.282$ $0.649$
$1\dots 50$ ($100\%$) $0.733$ $0.521$ $0.004$ $0.272$ $0.274$ $0.649$
(b) Deepfloyd IF
Appendix L Noise-to-image mapping
We showcase an additional study showing the differences that occur between noise and latent encodings, from the perspective of their mapping to the images. Several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. (2024) show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. (2024a), observing that even models with different architectures converge to the same score function and, hence, the same noise-to-imape mapping. Khrulkov and Oseledets (2022) show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works (Kim and Milman, 2011; Lavenant and Santambrogio, 2022) contradict this finding.
L.1 Smallest $l_{2}$ mapping
[…]
T ADM-32 ADM-64 ADM-256 LDM DiT
$4000$ $99.1_{\pm 1.0}$ $46.4_{\pm 3.0}$ $99.5_{\pm 0.3}$ $43.3_{\pm 6.7}$ - - - - - -
T ADM-32 ADM-64 ADM-256 LDM DiT
$4000$ $2.8_{\pm 2.2}$ $41.9_{\pm 3.0}$ $1.9_{\pm 1.4}$ $43.5_{\pm 6.5}$ - - - - - -
[…]
L.2 Asymmetry of Noise-to-Image mapping
Conclusion.
Appendix M On Noise-Image-Latent relations during diffusion training
To further explore the relationships that exist between noises, generations, and latents, we study how the relationships between them change with the training of the diffusion model. We train two diffusion models from scratch and follow the setup from Nichol and Dhariwal (2021) for two unconditional ADMs for the ImageNet ($64\times 64$) and CIFAR-10 ($32\times 32$) datasets. The CIFAR-10 model is trained for 700700K steps, while the ImageNet model – for 1.51.5M steps, both with a batch size of $128$. Models employ a cosine scheduler with 44K diffusion steps.
M.1 Spatial relations of noise and latents over training time
M.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from Appendix L over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In Fig. 16, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
(a) ADM-64
(b) ADM-32
M.3 Image alignment over training time
In Fig. 18, we visualize how the diffusion model learns the low-frequency features of the image already at the beginning of the training when comparing generations from the next training steps against the generations after finishing training for the ADM-32 model trained on the CIFAR-10 dataset. In  Fig. 19 (ADM-64) and Fig. 20 (ADM-32), we show additional examples illustrating how generations evolve over training for the same Gaussian noise $\mathbf{x_{T}}$} using a DDIM sampler. Initially, low-frequency features emerge and remain relatively stable, while continued training improves generation quality by refining only the high-frequency details.
Figure 18: Similarity of the generations sampled from the same random noise at different stages of the diffusion model’s training to the final outputs for ADM-32 (CIFAR-10). We plot CKA, SVCCA, SSIM, and DINO image alignment metrics and show that the diffusion model already learns the mapping between Gaussian noise and generations at the beginning of the training.
Appendix N Parameter impact analysis
N.1 Number of inversion steps